# Groupoid Graded Semisimple Rings

### v1

@@ 8. Semisimple categories
Analogously, we define the category $\operatorname{Fun_{gr}}(\mathcal{C},\,\varepsilon\mathcal{G}-\operatorname{mod\textrm{-}\!}\mathbb{Z}^{(\mathcal{G}_{0})})$ whose objects are the additive covariant functors $F:\mathcal{C}\to\varepsilon\mathcal{G}-\operatorname{mod\textrm{-}\!}\mathbb{Z}^{(\mathcal{G}_{0})}$ such that, if $X,Y\in\mathcal{C}_{0}$ and $f\in\mathcal{C}(X,Y)$, then $F(f)\in\operatorname{HOM}(F(X),F(Y))_{(Y,X)}$. Then, for a $\mathcal{G}$-graded left $R[\mathcal{C}]$-module $M$ we can build functors $M(\mathcal{C}_{0},-)$ and $M(A,-)$. In this way, we prove that there exists an additive covariant isomorphism of categories
$\mathcal{G}-R[\mathcal{C}]\operatorname{\!\textrm{-}gr}$ $\longrightarrow\operatorname{Fun_{gr}}(\mathcal{C},\,\varepsilon\mathcal{G}-\operatorname{mod\textrm{-}\!}\mathbb{Z}^{(\mathcal{G}_{0})})$
$M$ $\longmapsto M(\mathcal{C}_{0},-)$
whose inverse is given by $F\longmapsto M[F]$, where $M[F]_{(X,Y)}:=F(X)_{(X,Y)}$, for each $(X,Y)\in\mathcal{G}$. If $A\in\mathcal{C}_{0}$, then such isomorphism restricts to
@@
    an additive covariant isomorphism of categories
    $\mathcal{G}\varepsilon_{A}-R[\mathcal{C}]\operatorname{\!\textrm{-}gr}$ $\longrightarrow\operatorname{Fun}(\mathcal{C},\mathcal{A}b)$
    $M$ $\longmapsto M(A,-).$
@@
The two previous constructions are connected by the next theorem. Before stating the result, we need some definitions. Following [29, p. 18], we define an ideal of $\mathcal{C}$ as an additive subfunctor of the bifunctor $\mathcal{C}(-,-):\mathcal{C}^{op}\times\mathcal{C}\to\mathcal{A}b$. We denote the category of additive bifunctors $\mathcal{C}^{op}\times\mathcal{C}\to\mathcal{A}b$ by $\operatorname{Bifun}(\mathcal{C},\mathcal{A}b)$. We also define the category $R[\mathcal{C}]\operatorname{\!\textrm{-}gr\textrm{-}\!}R[\mathcal{C}]$ whose objects are the $\mathcal{G}$-graded $(R[\mathcal{C}],R[\mathcal{C}])$-bimodules and the morphisms are the homomorphisms of bimodules $g:M\to N$ such that $g(M_{\sigma})\subseteq N_{\sigma}$ for all $\sigma\in\Gamma$. Note that, by Remark 2.5(1), the second part of the following theorem contains a characterization of ideals of $R[\mathcal{C}]$.
@@
Let $\mathcal{G}:=\mathcal{C}_{0}\times\mathcal{C}_{0}$. There exists an additive covariant isomorphism of categories
$R[\mathcal{C}]\operatorname{\!\textrm{-}gr\textrm{-}\!}R[\mathcal{C}]$ $\longrightarrow\operatorname{Bifun}(\mathcal{C},\mathcal{A}b)$
$M$ $\longmapsto M(-,-).$
whose inverse is given by $F\longmapsto M[F]$. Such isomorphism restricts to a bijection
[…]
@@
The verification that $M\longmapsto M(-,-)$ and $F\longmapsto M[F]$ define mutually inverse functors is similar to the proof of Theorem 8.1.
For the second part of the statement, it suffices to note that if $M$ is a graded ideal of $R[\mathcal{C}]$, then $M(-,-)$ is a subfunctor of $\mathcal{C}(-,-)$ and, conversely, if $F$ is an ideal of $\mathcal{C}$, then $M[F]$ is a graded ideal of $R[\mathcal{C}]$. ∎
@@ 8.2. Semisimple categories
Following [29, p. 18], we will say that $\mathcal{C}$ is a (right) semisimple category if $\mathcal{C}(-,A)$ is a semisimple object in the category $\operatorname{Fun}(\mathcal{C}^{op},\mathcal{A}b)$ for all $A\in\mathcal{C}_{0}$.
@@
Let $\{\mathcal{C}_{j}:j\in J\}$ be a family of preadditive categories with at least a zero object. Following [13, p. 133], let $\prod_{j\in J}^{f}\mathcal{C}_{j}$ be the full subcategory of $\prod_{j\in J}\mathcal{C}_{j}$ whose objects are of the form $(A_{j})_{j\in J}$, where $A_{j}$ is a zero object of $\mathcal{C}_{j}$ for every $j\in J$ except for finitely many indices $j$.
We have the following characterization of semisimple categories. The equivalence of items (1) and (8) was first proved in [29, pp. 19-20].
@@
$(2)\implies(8):$ Suppose that $R[\mathcal{C}]$ is a gr-semisimple $\mathcal{G}$-graded ring. By Theorem 5.30, there exist a set $J$, a family $\{K_{j}\colon j\in J\}$ of nonempty sets, a sequence $(A_{j})_{j\in J}\in\mathcal{C}_{0}^{J}$, a summable family of $\mathcal{G}$-graded rings $\{\operatorname{M}_{K_{j}}(D_{j})(\overline{\varsigma}_{j})\colon j\in J\}$ and an gr-isomorphism of $\mathcal{G}$-graded rings
@@
$(8)\implies(9)$: It suffices to consider the functor $F:\mathcal{C}\to\operatorname{f.g.\!-mod\textrm{-}\!}\bigoplus\limits_{j\in J}D_{j}$ defined by $F(A)=\bigoplus_{j\in J}D_{j}^{n(j,A)}$.
@@
$(8)\implies(10)$: It suffices to consider the functor $F:\mathcal{C}\to\prod_{j\in J}^{f}\operatorname{f.g.\!-mod\textrm{-}\!}D_{j}$ defined by $F(A)=(D_{j}^{n(j,A)})_{j\in J}$.
@@
Let $R$ be a unital ring and suppose that $\mathcal{C}$ is a small full subcategory of $\operatorname{f.g.\!-mod\textrm{-}\!}R$ whose objects are semisimple. Given $M$ and $N$ objects of $\mathcal{C}$, we can write $M=\bigoplus_{i=1}^{m}S_{i}$ and $N=\bigoplus_{j=1}^{n}T_{j}$, where $S_{i}$ and $T_{j}$ are all simple $R$-modules. Then $\operatorname{Hom}_{R}(M,N)$ is identified with a finite product of finite matrix rings. By item (8) of Theorem 8.5, $\mathcal{C}$ is a semisimple category.∎
@@ 8.3. Simple artinian categories
Following [29, p. 18-19], we will say that $\mathcal{C}$ is a right artinian category if $\mathcal{C}(-,A)$ is an artinian object of the category $\operatorname{Fun}(\mathcal{C}^{op},\mathcal{A}b)$ for each $A\in\mathcal{C}_{0}$. That is, the subobjects of $\mathcal{C}(-,A)$ satisfy the descending chain condition. We will say that $\mathcal{C}$ is a simple category if the only nonzero ideal of $\mathcal{C}$ is $\mathcal{C}(-,-)$.
@@
It easy to see that $R[\mathcal{C}]$ is a gr-simple ring if and only if $\mathcal{C}$ is a simple category in the sense of [13, Section 7.2]. Therefore, it follows from Proposition 8.7(3) that our concept of simple category coincides with that of [13]. In [13, Section 7.2], it was proved that if $R$ is a simple (unital) ring, then the category $\operatorname{proj\textrm{-}\!}R$ of finitely generated projective right $R$-modules is a simple category [13, Proposition 7.6] and that a preadditive category $\mathcal{P}$ is a simple category if and only if there exists a simple ring $R$ such that $\mathcal{P}$ is equivalent to a full subcategory of $\operatorname{proj\textrm{-}\!}R$ [13, Theorem 7.5]. The following result characterizes the simple artinian categories.
@@
$(4)\implies(5)$: It suffices to consider the functor $F:\mathcal{C}\to\operatorname{f.g.\!-mod\textrm{-}\!}D$ defined by $F(A)=D^{n(A)}$.
@@ 8.4. Free functors
In this subsection, we explore when $R[\mathcal{C}]$, the ring of the small preadditive category $\mathcal{C}$, is a pfm ring and relate this fact with the concept of free functors defined in [29, p. 17-18]. We proceed to recall the definition of that notion. Let $F:\mathcal{C}\to\mathcal{A}b$ be an additive contravariant functor. We say that $(x_{i})_{i\in I}\in\prod_{i\in I}F(A_{i})$, where $(A_{i})_{i\in I}\in\mathcal{C}_{0}^{I}$, is a sequence of generators of $F$ if, for all $A\in\mathcal{C}_{0}$ and $x\in F(A)$, there exists $(\lambda_{i})_{i\in I}\in\bigoplus_{i\in I}\mathcal{C}(A,A_{i})$ such that
@@
The next result associates free functors with pseudo-free $R[\mathcal{C}]$-modules. The equivalence of (1) and (3) was given in [29, p. 18].
@@
    There exists a family :=$\mathcal{F}:=\{D_{j}:j\in J\}$ of division rings such that $\mathcal{C}$ is isomorphic to a small full subcategory of $\operatorname{f.g.\!-mod\textrm{-}\!}\bigoplus\limits_{j\in J}D_{j}$ whose set of objects contains $\mathcal{F}$.
@@
    There exists a family :=$\mathcal{F}:=\{D_{j}:j\in J\}$ of division rings such that $\mathcal{C}$ is isomorphic to a small full subcategory of $\prod_{j\in J}^{f}\operatorname{f.g.\!-mod\textrm{-}\!}D_{j}$ whose set of objects contains $\mathcal{F}$.
@@
Let $A$ be a ring with unity. Consider a small preadditive subcategory $\mathcal{A}$ of the finitely generated semisimple right $A$-modules that contains a set of representatives $\{S_{j}\colon j\in J\}$ of simple right $A$-modules. Then $R[\mathcal{A}]$, the ring of the category $\mathcal{A}$, satisfies Theorem 8.10(4). Thus, $R[\mathcal{A}]$ is a pfm ring. Therefore, suppose that $\mathcal{B}$ is a small preadditive category such that $R[\mathcal{B}]$ is not pfm. Then $\mathcal{B}$ cannot be equivalent to the category of finitely generated semisimple modules of a ring.
Categories of finitely generated semisimple modules of a ring are examples of amenable semisimple categories [29, p. 20]. The latter categories were characterized in [13, Theorem 4.55]. We note that small amenable semisimple categories are examples of categories satisfying Theorem 8.10.
@@ 8.5. Division categories
$(2)\iff(3)$: By Theorem 7.6, (2) holds if, and only if, $R[\mathcal{C}]$ is a pfm $\mathcal{G}$-graded ring and, for each $\mathcal{G}$-graded $R[\mathcal{C}]$-module $M$ all pseudo-basis of $M$ have the same cardinality. By Theorem 8.10 and Proposition 8.9, this is equivalent to (3).
$(2)\iff(4)$: By Theorem 7.8, (2) holds if, and only if, $R[\mathcal{C}]$ is a gr-semisimple $\mathcal{G}$-graded ring and $R[\mathcal{C}]_{(A,A)}$ is a division ring for all nonzero $A\in\mathcal{C}_{0}$. By Theorem 8.5, this is equivalent to $\mathcal{C}$ being a semisimple category and $\mathcal{C}(A,A)$ being a division ring for all $A$ such that $\mathcal{C}(A,A)\neq 0$.

### v2

@@ 8. Semisimple categories
Analogously, we define the category $\operatorname{Fun_{gr}}(\mathcal{C},\,\varepsilon\mathcal{G}-\operatorname{mod\textrm{-}\!}\mathbb{Z}^{(\mathcal{G}_{0})})$ whose objects are the additive covariant functors $F:\mathcal{C}\to\varepsilon\mathcal{G}-\operatorname{mod\textrm{-}\!}\mathbb{Z}^{(\mathcal{G}_{0})}$ such that, if $X,Y\in\mathcal{C}_{0}$ and $f\in\mathcal{C}(X,Y)$, then $F(f)\in\operatorname{HOM}(F(X),F(Y))_{(Y,X)}$. Then, for a $\mathcal{G}$-graded left $R[\mathcal{C}]$-module $M$ we can build functors $M(\mathcal{C}_{0},-)$ and M(A,)M(A,-). In this way, we prove that there exists an equivalence of categories
$\mathcal{G}-R[\mathcal{C}]\operatorname{\!\textrm{-}gr}$ $\leftrightarrows\operatorname{Fun_{gr}}(\mathcal{C},\,\varepsilon\mathcal{G}-\operatorname{mod\textrm{-}\!}\mathbb{Z}^{(\mathcal{G}_{0})})$
$M$ $\mapsto M(\mathcal{C}_{0},-)$
$M[F]$ $\mapsfrom F,$
where $M[F]_{(X,Y)}:=F(X)_{(X,Y)}$ for each $(X,Y)\in\mathcal{G}$. If $A\in\mathcal{C}_{0}$, then we get
@@
    an equivalence of categories
    $\mathcal{G}\varepsilon_{A}-R[\mathcal{C}]\operatorname{\!\textrm{-}gr}$ $\rightarrow\operatorname{Fun}(\mathcal{C},\mathcal{A}b)$
    $M$ $\mapsto M(A,-).$
@@
The next theorem links the two previous constructions. Before stating the result, we need some definitions. Following [32, p. 18], we define an ideal of $\mathcal{C}$ as an additive subfunctor of the bifunctor $\mathcal{C}(-,-):\mathcal{C}^{op}\times\mathcal{C}\to\mathcal{A}b$. We denote the category of additive bifunctors $\mathcal{C}^{op}\times\mathcal{C}\to\mathcal{A}b$ by $\operatorname{Bifun}(\mathcal{C},\mathcal{A}b)$. We also define the category $R[\mathcal{C}]\operatorname{\!\textrm{-}gr\textrm{-}\!}R[\mathcal{C}]$ whose objects are the $\mathcal{G}$-graded $(R[\mathcal{C}],R[\mathcal{C}])$-bimodules and the morphisms are the homomorphisms of bimodules $g:M\to N$ such that $g(M_{\sigma})\subseteq N_{\sigma}$ for all $\sigma\in\Gamma$. Note that, by Remark 2.5(1), the second part of the following theorem contains a characterization of ideals of $R[\mathcal{C}]$.
@@
Let $\mathcal{G}:=\mathcal{C}_{0}\times\mathcal{C}_{0}$. The following functors define an equivalence of categories
$R[\mathcal{C}]\operatorname{\!\textrm{-}gr\textrm{-}\!}R[\mathcal{C}]$ $\leftrightarrows\operatorname{Bifun}(\mathcal{C},\mathcal{A}b)$
$M$ $\mapsto M(-,-)$
$M[F]$ $\mapsfrom F;$
Moreover, such equivalence induces a bijection
$\{\textrm{graded ideals of $R[\mathcal{C}]$}\}\longrightarrow\{\textrm{ideals of $\mathcal{C}$}\}$
@@
The verification that $M\mapsto M(-,-)$ and $F\mapsto M[F]$ define an equivalence is similar to the proof of Theorem 8.1.
For the second part of the statement, it suffices to note that if $M$ is a graded ideal of $R[\mathcal{C}]$, then M(,)M(-,-) is a subfunctor of $\mathcal{C}(-,-)$ and, conversely, if $F$ is an ideal of $\mathcal{C}$, then $M[F]$ is naturally gr-isomorphic to a graded ideal of $R[\mathcal{C}]$. ∎
@@ 8.2. Semisimple categories
Following [32, p. 18], we will say that $\mathcal{C}$ is a (right) semisimple category if $\mathcal{C}(-,A)$ is a semisimple object in the category $\operatorname{Fun}(\mathcal{C}^{op},\mathcal{A}b)$ for all $A\in\mathcal{C}_{0}$.
@@
Let $\{\mathcal{C}_{j}:j\in J\}$ be a family of preadditive categories with at least a zero object. Following [16, p. 133], let $\prod_{j\in J}^{f}\mathcal{C}_{j}$ be the full subcategory of $\prod_{j\in J}\mathcal{C}_{j}$ whose objects are of the form $(A_{j})_{j\in J}$, where $A_{j}$ is a zero object of $\mathcal{C}_{j}$ for every $j\in J$ except for finitely many indices $j$.
We have the following characterization of semisimple categories. The equivalence of items (1) and (8) was first proved in [32, pp. 19-20].
@@
$(2)\implies(8):$ Suppose that $R[\mathcal{C}]$ is a gr-semisimple $\mathcal{G}$-graded ring. By Theorem 5.30, there exist a set $J$, a family $\{K_{j}\colon j\in J\}$ of non-empty sets, a sequence $(A_{j})_{j\in J}\in\mathcal{C}_{0}^{J}$, a summable family of $\mathcal{G}$-graded rings $\{\operatorname{M}_{K_{j}}(D_{j})(\overline{\varsigma}_{j})\colon j\in J\}$ and an gr-isomorphism of $\mathcal{G}$-graded rings
@@
$(8)\implies(9)$: It suffices to consider the functor $F:\mathcal{C}\to\operatorname{f.g.\!-mod\textrm{-}\!}\bigoplus\limits_{j\in J}D_{j}$ defined on objects by $F(A)=\bigoplus_{j\in J}D_{j}^{n(j,A)}$ and defined on morphisms using the maps $\varphi_{(A,B)}$.
@@
$(8)\implies(10)$: It suffices to consider the functor $F:\mathcal{C}\to\prod_{j\in J}^{f}\operatorname{f.g.\!-mod\textrm{-}\!}D_{j}$ defined on objects by $F(A)=(D_{j}^{n(j,A)})_{j\in J}$, and on morphisms using the maps $\varphi_{(A,B)}$.
@@
Let $R$ be a unital ring and let $\{S_{j}:j\in J\}$ be a complete set of representatives of the isomorphism classes of simple right $R$-modules. Set $D_{j}:=\operatorname{Hom}_{R}(S_{j},S_{j})$ for all $j\in J$.
Suppose that $\mathcal{C}$ is a small full subcategory of $\operatorname{f.g.\!-mod\textrm{-}\!}R$ whose objects are semisimple modules. Given objects $M$ and $N$ of $\mathcal{C}$, there exist sequences $(n(j,M))_{j\in J}$, $(n(j,N))_{j\in J}\in\mathbb{N}^{J}$ such that
$M\cong\bigoplus_{j\in J}S_{j}^{n(j,M)},\quad\quad N\cong\bigoplus_{j\in J}S_{j}^{n(j,N)},$
$\sum\limits_{j\in J}n(j,M)<\infty$ and $\sum\limits_{j\in J}n(j,N)<\infty$. Then $\operatorname{Hom}_{R}(M,N)$ is naturally identified with
$\prod_{j\in J}\operatorname{M}_{n(j,N)\times n(j,M)}(D_{j})$
as in Theorem 8.5(8). Hence, $\mathcal{C}$ is a semisimple category. ∎
@@ 8.3. Simple artinian categories
Following [32, p. 18-19], we will say that $\mathcal{C}$ is a right artinian category if $\mathcal{C}(-,A)$ is an artinian object of the category $\operatorname{Fun}(\mathcal{C}^{op},\mathcal{A}b)$ for each $A\in\mathcal{C}_{0}$. That is, the subobjects of $\mathcal{C}(-,A)$ satisfy the descending chain condition. We will say that $\mathcal{C}$ is a simple category if the only nonzero ideal of $\mathcal{C}$ is $\mathcal{C}(-,-)$.
@@
It is easy to see that $R[\mathcal{C}]$ is a gr-simple ring if and only if $\mathcal{C}$ is a simple category in the sense of [16, Section 7.2]. Therefore, it follows from Proposition 8.7(3) that our concept of simple category coincides with that of [16]. In [16, Section 7.2], it was proved that if $R$ is a simple (unital) ring, then the category $\operatorname{proj\textrm{-}\!}R$ of finitely generated projective right $R$-modules is a simple category [16, Proposition 7.6] and that a preadditive category $\mathcal{P}$ is a simple category if and only if there exists a simple ring $R$ such that $\mathcal{P}$ is equivalent to a full subcategory of $\operatorname{proj\textrm{-}\!}R$ [16, Theorem 7.5]. The following result characterizes the simple artinian categories.
@@
$(4)\implies(5)$: It suffices to consider the functor $F:\mathcal{C}\to\operatorname{f.g.\!-mod\textrm{-}\!}D$ defined on objects by $F(A)=D^{n(A)}$, and defined on morphisms using the maps $\varphi_{(A,B)}$.
@@ 8.4. Free functors
In this subsection, we explore when $R[\mathcal{C}]$, the ring of the small preadditive category $\mathcal{C}$, is a pfm ring and relate this fact with the concept of free functors defined in [32, p. 17-18]. We proceed to recall the definition of that notion. Let $F:\mathcal{C}\to\mathcal{A}b$ be an additive contravariant functor. We say that $(x_{i})_{i\in I}\in\prod_{i\in I}F(A_{i})$, where $(A_{i})_{i\in I}\in\mathcal{C}_{0}^{I}$, is a sequence of generators of $F$ if, for all $A\in\mathcal{C}_{0}$ and $x\in F(A)$, there exists $(\lambda_{i})_{i\in I}\in\bigoplus_{i\in I}\mathcal{C}(A,A_{i})$ such that
@@
The next result associates free functors with pseudo-free $R[\mathcal{C}]$-modules. The equivalence of (1) and (3) was given in [32, p. 18].
@@
    There exists a family $\mathcal{F}:=\{D_{j}:j\in J\}$ of division rings such that $\mathcal{C}$ is isomorphic to a small full subcategory of $\operatorname{f.g.\!-mod\textrm{-}\!}\bigoplus\limits_{j\in J}D_{j}$ whose set of objects contains $\mathcal{F}$.
@@
    There exists a family $\mathcal{F}:=\{D_{j}:j\in J\}$ of division rings such that $\mathcal{C}$ is isomorphic to a small full subcategory of $\prod_{j\in J}^{f}\operatorname{f.g.\!-mod\textrm{-}\!}D_{j}$ whose set of objects contains $\mathcal{F}$.
@@
Let $R$ be a unital ring and let $\mathcal{C}$ be a small full subcategory of $\operatorname{f.g.\!-mod\textrm{-}\!}R$. Suppose that there exists a subset $\mathcal{S}\subseteq\mathcal{C}_{0}$ of simple right $R$-modules such that all the objects of $\mathcal{C}$ are isomorphic to finite direct sums of elements of $\mathcal{S}$. Then $R[\mathcal{C}]$, the ring of the category $\mathcal{C}$, satisfies Theorem 8.10(4). Thus, $R[\mathcal{C}]$ is a pfm ring. Therefore, suppose that $\mathcal{B}$ is a small preadditive category such that $R[\mathcal{B}]$ is not pfm. Then $\mathcal{B}$ cannot be equivalent to the category of finitely generated semisimple modules of a ring.
The category of all finitely generated semisimple modules of a ring is an example of an amenable semisimple category [32, p. 20]. A characterization of such categories is given in [16, Theorem 4.55]. We observe that small amenable semisimple categories satisfy the conditions of Theorem 8.10.
@@ 8.5. Division categories
$(2)\iff(3)$: By Theorem 7.6, (2) holds if and only if $R[\mathcal{C}]$ is a pfm $\mathcal{G}$-graded ring and, for each $\mathcal{G}$-graded $R[\mathcal{C}]$-module $M$ all pseudo-basis of $M$ have the same cardinality. By Theorem 8.10 and Proposition 8.9, this is equivalent to (3).
$(2)\iff(4)$: By Theorem 7.8, (2) holds if and only if $R[\mathcal{C}]$ is a gr-semisimple $\mathcal{G}$-graded ring and $R[\mathcal{C}]_{(A,A)}$ is a division ring for all nonzero $A\in\mathcal{C}_{0}$. By Theorem 8.5, this is equivalent to $\mathcal{C}$ being a semisimple category and $\mathcal{C}(A,A)$ being a division ring for all $A$ such that $\mathcal{C}(A,A)\neq 0$.
@@
Acknowledgments
We thank the referee for helpful suggestions and corrections that improved the exposition.